OpenStudy (anonymous):

The height of an object tossed upward with an initial velocity of 152 feet per second is given by the formula h = −16t2 + 152t, where h is the height in feet and t is the time in seconds. Find the time required for the object to return to its point of departure.

OpenStudy (anonymous):

first u need to find the the highest point where it has final speed zero... then u have to calulate its time ... after that u have to find the time when the hight becomes zero while it hits the ground .. so add these two times u will have your answer

OpenStudy (anonymous):

and u have to assume the innitial hight to be zero or ground reference

OpenStudy (anonymous):

Initial height is zero, so set the function equal to zero and solve.\[h(t) = −16t^2 + 152t=0 \implies t(t-\frac{19}{2})=0 \implies t=0~~~or~~~t=\frac{19}{2}\]At t=0 the object is launched; at t=19/2 it lands.